# Out-of-sample gravity predictions

### v1

@@ Abstract
Gravity equations are often used to evaluate counterfactual trade policy scenarios, such as the effect of regional trade agreements on trade flows. In this paper, we argue that the suitability of gravity equations for this purpose crucially depends on their out-of-sample predictive power. We propose a methodology that compares different versions of the gravity equation, both among themselves and with machine learning-based forecast methods such as random forests and neural networks. We find that the 3-way gravity model is difficult to beat in terms of out-of-sample average predictive performance, further justifying its place as the predominant tool for applied trade policy analysis. However, when the goal is to predict individual bilateral trade flows, the 3-way model can be outperformed by an ensemble machine learning method.
@@ 1 Introduction
Gravity equations are by far the most popular tool in international trade for evaluating the effects of policy interventions such as regional trade agreements (RTAs). In its most basic form, the gravity model states that exports from country $i$ to country $j$ at time $t$ are a log-linear function of exporter and importer characteristics (e.g., GDP) and bilateral variables such as distance.
@@
We believe that providing evidence on the relative out-of-sample predictive performance of the gravity equation is important for two reasons. [^4] Note that we will be using the terms predicting and forecasting interchangeably throughout this paper. First, having high-quality statistical models for predicting trade flows is directly relevant for policymaking. This often requires making such forecasts, for example to predict the potential trade flows for a given pair of countries. [^5] For example, the International Trade Centre has developed a so-called ‘export potential assessment methodology’ in order to help countries identify promising products for export promotion activities (see ITC, 2025). The more accurate the forecast of potential trade flows between two countries, the better such export promotion activities can be targeted.
@@
Our results suggest that the gravity equation performs well against these highly flexible forecasting tools, suggesting that it captures essential features of the data generating process of international trade flows. However, this is only true for the 3-way gravity model, which controls for exporter-time, importer-time, and bilateral fixed effects. By contrast, the traditional gravity equation (no fixed effects) and the 2-way gravity model (exporter-time, importer-time fixed effects only) can be outperformed by some of the machine learning methods we consider. This underlines the importance of bilateral fixed effects in the 3-way gravity model. Given that the 3-way model is already considered to be the state-of-the-art version of the gravity equation (see Yotov et al., 2016), we see our results as a further justification for its use.
@@ 2 Gravity’s out-of-sample forecast performance and causal estimates of trade policy interventions
where the summation in both the numerator and denominator is across all observations with $D=1$, i.e., the “treated” observations. Furthermore, $\hat{y}_{ijt}^{0}$ is the predicted value of $y_{ijt}^{0}$ based on the model and estimator being used. If the sum in the denominator $\sum_{D=1}\hat{y}_{ijt}^{0}$ consistently estimates $\sum_{D=1}y_{ijt}^{0}$, expression (1) gives a consistent estimate for the multiplicative treatment effect $e^{\tau}$. Along the same lines, less biased predictions for each $y_{ijt}^{0}$ will result in less biased estimates of the treatment effect, and predictions with less variance will result in increased precision.
@@ 3 Evaluating out-of-sample predictions
    Split the relevant dataset into a training and a test sample; we describe how we split the sample in Section 5.3.
@@ 4 Statistical models for trade flows
Applying machine learning estimators to our data requires choosing the set of predictors to be used. We consider two cases. We start by using just the covariates used by the traditional gravity equation (e.g., GDPs and distance). Additionally, we also include as regressors the three sets of fixed effects estimated from the 3-way gravity model. [^8] The alternative of including the full set of dummies representing the fixed effects is computationally infeasible. Including the fixed effects ensures that the machine learning methods have access to the same information as the 3-way specification, enabling a fairer comparison between the parametric and nonparametric approaches.
@@
Finally, we estimate a “3-way-ML” gravity specification, which augments the 3-way gravity model with the logarithm of the fitted values from the ensemble without fixed effects. This specification allows us to assess whether the parametric 3-way model can be improved by incorporating flexible, nonlinear predictions from the machine learning methods. [^9] Note that the ensemble can predict zero trade flows, which makes it impossible to include the log of those forecast values. To address this, we impose that the 3-way-ML model also predict zero trade whenever the ensemble without fixed effects does so. We acknowledge that this modeling choice is ad hoc and may limit the effectiveness of integrating machine learning predictions into the gravity framework.
@@ 5.1 Measures of predictive accuracy
We discuss in more detail below how we select the observations to be predicted, but for now simply note that there will be $n$ such observations. These $n$ observations correspond to our test sample in machine learning terminology, with the remaining NnN-n observations in our data serving as the estimation (or training) sample, where $N$ is the total number of observations in our dataset. We repeat the procedure of splitting the data into test and estimation sample and calculating our measures of predictive accuracy $K=1,000$ times. We then compute averages of our measures across the $K$ repetitions.
@@ 5.2 Data
We implement our tests of the out-of-sample performance of the estimators described in Section 4 on a standard and publicly available dataset of bilateral trade flows, that of Yotov (2025), which in turn is based on the United Nation’s Comtrade database. [^11] These data can be downloaded here: https://yotoyotov.com/Gravity_Undergrads.html The dataset includes aggregate trade and gravity variables for the $100$ largest exporters in the world, covering $98.9\%$ of world exports, $97.7\%$ of world imports, and $98.3\%$ of world GDP between 1990-2023.
@@ 5.3 Simulation design
    Second, we estimate a 3-way gravity model using the ppmlhdfe command of Correia et al. (2020), which implements Gourieroux et al.’s (1984) PPML estimator, as recommended by Santos Silva and Tenreyro (2006). The specification includes time-varying indicators for whether both countries in a pair are members of the European Union, whether they share a customs union, whether they have a regional trade agreement (RTA), and whether the importer has imposed a trade sanction on the exporter; these regressors will also be used in all subsequent regressions. The model also includes exporter-time, importer-time and bilateral fixed effects, whose estimated values are saved.
@@ 5.4 Results
RMAE 0.562 0.477 0.318 0.260 0.432 0.432 0.566 0.526 0.243 0.258 0.425 0.263 0.404
RRMSE 2.360 2.310 1.840 1.559 2.348 2.348 2.448 2.245 1.026 1.217 1.554 1.280 2.395
$R^{2}$ 0.718 0.781 0.890 0.914 0.791 0.791 0.718 0.747 0.913 0.906 0.817 0.913 0.807
Notes: This table shows simulation results for the case of endogenous treatment. The “Estimation” part focuses on the goal of obtaining consistent estimates of the effects of policy interventions (average forecasts), whereas the “Prediction” part focuses on the accuracy of forecasts for individual observations. The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way). The remaining columns refer to machine learning methods. See the text for details.
@@
Moving on to the machine learning methods, we find that some methods can do as well, or even better, than the 3-way model in terms of the mean of $IE_{k}$ (Ens, RF, GB-FE, and NN-FE), but have much higher $SE_{IE}$ (see the estimation part of the table). In turn, E-FE (the ensemble method augmented with the estimated fixed effects from the 3-way model) has the smallest $SE_{IE}$ and a mean of $IE_{k}$ reasonably close to $1$. As a result, the 3-way and E-FE estimators have virtually identical values of the $MSE_{IE}$ ($0.009$) and dominate all other methods according to this criterion.
@@
Finally, we note that the 3-way-ML estimator performs worse than the simpler 3-way gravity equation. Although this deserves further scrutiny, we conjecture that the deterioration in performance is due to overfitting. In any case, this is a reassuring result in the sense that it suggests that there is little reason to consider models more complex than the 3-way gravity equation.
We thus conclude that if a researcher is interested in obtaining consistent estimates of trade policy interventions such as RTAs, the 3-way gravity model is hard to beat. It is only when the researcher is interested in predicting individual trade flows out-of-sample that machine learning techniques can improve upon the performance of the 3-way model. However, this improvement is only possible if these techniques take account of the information from the fixed effects; the machine learning models not making use of these fixed estimates do not offer such an improvement. [^12] However, they do often outperform the traditional gravity equation that uses the same information set.
@@
RMAE 0.639 0.519 0.330 0.268 0.491 0.491 0.670 0.602 0.255 0.270 0.477 0.274 0.440
RRMSE 3.348 3.299 2.565 2.158 3.435 3.435 3.480 3.221 1.460 1.711 2.152 1.756 3.526
$R^{2}$ 0.701 0.782 0.882 0.906 0.773 0.773 0.690 0.720 0.903 0.896 0.795 0.904 0.800
Notes: This table shows simulation results for the case of exogenous treatment. The “Estimation” part focuses on the goal of obtaining consistent estimates of the effects of policy interventions (average forecasts), whereas the “Prediction” part focuses on the accuracy of forecasts for individual observations. The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way). The remaining columns refer to machine learning methods. See the text for details.
The results from Table 2 largely confirm those from Table 1.131313We also consider a case with $a=7.5$ and $b=1$, which leads to values of $n_{k}$ between $10$ and $203$, with a mean of $84.2$, and the results are qualitatively similar to those in Table 1. The 3-way model now outperforms all competitors in terms of the mean of the imputation estimator. However, the E-FE estimator beats it in terms of $SE_{IE}$, and the two estimators have again virtually indistinguishable values of $MSE_{IE}$. The one notable difference to Table 1 is the better performance of the 2-way model, which is now comparable to the 3-way model in terms of the bias of the imputation estimator and outperforms both the traditional gravity model and the 1-way model in this respect. This confirms our earlier supposition that the endogeneity of the treatment explains the poor performance of the 2-way model in Table 1. Turning now to the bottom part of the table, the E-FE approach is again the best performing method, both for the RMAE and RRMSE but not for the out-of-sample $R^{2}$; the 3-way model outperforms all other machine learning methods in terms of the RMAE and out-of-sample $R^{2}$.
@@ 6 Conclusion
We find that the 3-way gravity model is difficult to beat when the purpose is to obtain consistent estimates of the effects of policy interventions. The only area where machine learning methods can outperform the 3-way model is the prediction of individual trade flows. Given that the evaluation of policy interventions is arguably the most important reason for the use of the gravity equation in applied international trade research, we see this result as reassuring, in the sense that it further cements the role of the 3-way gravity model as the state-of-the-art specification for gravity estimation. Future research will hopefully adopt our methodological approach and investigate whether our results hold in different datasets and using different machine learning estimators.

### v2

@@ Abstract
Gravity equations are often used to evaluate counterfactual trade policy scenarios, such as the effect of regional trade agreements on trade flows. In this paper, we argue that the suitability of gravity equations for this purpose crucially depends on their out-of-sample predictive power. We propose a methodology that compares different versions of the gravity equation, both among themselves and with machine learning-based forecast methods such as random forests and neural networks. We find that the 3-way gravity model is difficult to beat in terms of out-of-sample average predictive performance, especially if a flexible specification is used. This result further justifies its place as the predominant tool for applied trade policy analysis. However, when the goal is to predict individual bilateral trade flows, the 3-way model can be outperformed by an ensemble machine learning method.
@@ 1 Introduction
Gravity equations are by far the most popular tool in international trade for evaluating the effects of policy interventions such as regional trade agreements (RTAs). In its most basic form, the gravity model states that exports from country $i$ to country $j$ at time $t$ are a log-linear function of exporter and importer characteristics (e.g., GDP) and bilateral variables (e.g., distance).
@@
We believe that providing evidence on the relative out-of-sample predictive performance of the gravity equation is important for two reasons. [^4] Note that we will be using the terms predicting and forecasting interchangeably throughout this paper. First, having high-quality statistical models for predicting trade flows is directly relevant for policymaking. This often requires making such forecasts, for example to predict the potential trade flows for a given pair of countries. For example, the International Trade Centre has developed a so-called ‘export potential assessment methodology’ in order to help countries identify promising products for export promotion activities (see ITC, 2025). The more accurate the forecast of potential trade flows between two countries, the better such export promotion activities can be targeted.
@@
Our results suggest that the gravity equation performs well against these highly flexible forecasting tools, suggesting that it captures essential features of the data generating process of international trade flows. However, this is only true for the 3-way gravity model, which controls for exporter-time, importer-time, and bilateral fixed effects. By contrast, the traditional gravity equation (no fixed effects) and the 2-way gravity model (exporter-time, importer-time fixed effects only) can be outperformed by some of the machine learning methods we consider. This underlines the importance of bilateral fixed effects in the 3-way gravity model. Given that the 3-way model is already considered to be the state-of-the-art version of the gravity equation (see Yotov et al., 2016), we see our results as a further justification for its use. However, our results also suggest that there may be (modest) gains from using a flexible specification of the way pair characteristics with time variation enter the model.
@@ 2 Gravity’s out-of-sample forecast performance and causal estimates of trade policy interventions
where the summation in both the numerator and denominator is across all observations with $D=1$, i.e., the “treated” observations. Furthermore, $\hat{y}_{ijt}^{0}$ is the predicted value of $y_{ijt}^{0}$ obtained using only the observations with $D=0$ and the selected model and estimator. [^5] Notice that the standard PPML estimator of $e^{\tau}$ is similar to (1), except that the values in the denominator are estimated using the full sample. If the sum in the denominator $\sum_{D=1}\hat{y}_{ijt}^{0}$ consistently estimates $\sum_{D=1}y_{ijt}^{0}$, expression (1) gives a consistent estimate for the multiplicative treatment effect $e^{\tau}$. Along the same lines, less biased predictions for each $y_{ijt}^{0}$ will result in less biased estimates of the treatment effect, and predictions with less variance will result in increased precision.
@@ 3 Evaluating out-of-sample predictions
    Split the relevant dataset into a training and a test sample; we describe how we split the data in Section 5.3.
@@ 4 Statistical models for trade flows
Applying machine learning estimators to our data requires choosing the set of predictors to be used. We consider two cases. We start by using just the covariates used by the traditional gravity equation (e.g., GDPs and distance). Additionally, we also include as regressors the three sets of fixed effects estimated with the 3-way gravity model for the corresponding training sample. [^8] The alternative of including the full set of dummies representing the fixed effects is computationally infeasible. Including the fixed effects ensures that the machine learning methods have access to the same information as the 3-way specification, enabling a fairer comparison between the parametric and nonparametric approaches.
@@
Finally, we estimate a “3-way-ML” gravity specification, which augments the 3-way gravity model with the logarithm of the fitted values from the neural network estimator without fixed effects, which are necessarily positive. This specification allows us to assess whether the parametric 3-way model can be improved by incorporating flexible, nonlinear predictions from the machine learning methods. [^9] Note that the predictions from the other machine learning methods are not necessarily positive, and therefore cannot be used to augment the 3-way model in the same way.
@@ 5.1 Measures of predictive accuracy
We discuss in more detail below how we select the observations to be predicted, but for now simply note that there will be $n$ such observations. These $n$ observations correspond to our test sample in machine learning terminology, with the remaining NnN-n observations in our data serving as the estimation (or training) sample, where $N$ is the total number of observations in our dataset. We repeat the procedure of splitting the data into test and estimation sample and calculating our measures of predictive accuracy $K=1,000$ times. We then compute descriptive statistics of our measures across the $K$ repetitions.
@@ 5.2 Data
We implement our tests of the out-of-sample performance of the estimators described in Section 4 on a standard and publicly available dataset of bilateral trade flows, that of Yotov (2025), which in turn is based on the United Nation’s Comtrade database. [^11] These data can be downloaded here: https://yotoyotov.com/Gravity_Undergrads.html The dataset includes aggregate trade and gravity variables for the $100$ largest exporters in the world, covering $98.9\%$ of world exports, $97.7\%$ of world imports, and $98.3\%$ of world GDP between 1990-2023. Besides data on trade, the database includes standard explanatory variables that include a measure of distance, GDPs for importer and exporter, and indicators for membership of the European Union, membership of a customs union, regional trade agreements, contiguity, common language, and colonial ties, as well as an indicator for whether the importer has imposed a trade sanction on the exporter. All these regressors are used in all subsequent regressions, although in come cases they drop out due to collinearity with the fixed effects.
@@ 5.3 Simulation design
    Second, we estimate a 3-way gravity model using the ppmlhdfe command of Correia et al. (2020), which implements Gourieroux et al.’s (1984) PPML estimator, as recommended by Santos Silva and Tenreyro (2006). The specification includes all the regressors described above, as well as exporter-time, importer-time and bilateral fixed effects, whose estimated values are saved.
@@ 5.4 Results
RMAE 0.562 0.477 0.318 0.260 0.432 0.432 0.566 0.526 0.243 0.258 0.425 0.263 0.244
RRMSE 2.360 2.310 1.840 1.559 2.348 2.348 2.448 2.245 1.026 1.217 1.554 1.280 1.114
$R^{2}$ 0.718 0.781 0.890 0.914 0.791 0.791 0.718 0.747 0.913 0.906 0.817 0.913 0.917
Notes: This table shows simulation results for the case of endogenous treatment. The “Estimation” part focuses on the goal of obtaining consistent estimates of the effects of policy interventions (average forecasts), whereas the “Prediction” part focuses on the accuracy of forecasts for individual observations. The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way). The remaining columns refer to machine-learning based methods. See the text for details.
@@
Moving on to the machine learning methods, we find that some methods can do as well, or even better, than the 3-way model in terms of the mean of $IE_{k}$ (Ens, RF, GB-FE, and NN-FE), but have much higher $SE_{IE}$ (see the estimation part of the table). In turn, E-FE (the ensemble method augmented with the estimated fixed effects from the 3-way model) has a smaller $SE_{IE}$ than the 3-way model and a mean of $IE_{k}$ reasonably close to $1$. As a result, the 3-way and E-FE estimators have virtually identical values of the $MSE_{IE}$ ($0.009$).
@@
Finally, we find that the 3-way-ML estimator outperforms all other estimators we considered in terms of estimation criteria, and has the best out-of-sample $R^{2}$. Since the model includes 3-way fixed effects, the improvement over the basic 3-way model can only come from the augmented model that better captures the effect of pair characteristics with time variation, such as regional trade agreements and customs unions. This result is not surprising because it is well known that these can have very heterogeneous effects (see, e.g., Baier et al., 2019), and we model them with single dummies. Therefore, our results suggest that, when the aim is to evaluate policy interventions, it may be useful to employ a flexible specification of the 3-way model that more directly captures the effect of time-varying pair characteristics. This can be done parametrically (for example, by including interactions of the standard dummies with partners’ characteristics), or nonparametrically as we do in the 3-way-ML estimator. [^12] The advantage of the nonparametric way of controlling for these variables is that it avoids the potential overfitting from rich parametric specifications. In any case, our results suggest that the gains from using more complex models is relatively limited.
We thus conclude that if a researcher is interested in obtaining consistent estimates of trade policy interventions such as RTAs, the 3-way gravity model is hard to beat, especially if it is flexibly specified. It is only when the researcher is interested in predicting individual trade flows out-of-sample that machine learning techniques can improve upon the performance of the 3-way models. However, this improvement is marginal and is only possible if these techniques take account of the information from the fixed effects; the machine learning models not making use of these fixed estimates do not offer such an improvement. [^13] However, they often outperform the traditional gravity equation that uses the same information set.
@@
RMAE 0.639 0.519 0.330 0.268 0.491 0.491 0.670 0.602 0.255 0.270 0.477 0.274 0.259
RRMSE 3.348 3.299 2.565 2.158 3.435 3.435 3.480 3.221 1.460 1.711 2.152 1.756 1.742
$R^{2}$ 0.701 0.782 0.882 0.906 0.773 0.773 0.690 0.720 0.903 0.896 0.795 0.904 0.908
Notes: This table shows simulation results for the case of exogenous treatment. The “Estimation” part focuses on the goal of obtaining consistent estimates of the effects of policy interventions (average forecasts), whereas the “Prediction” part focuses on the accuracy of forecasts for individual observations. The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way). The remaining columns refer to machine-learning based methods. See the text for details.
The results from Table 2 largely confirm those from Table 1.141414We also consider a case with $a=7.5$ and $b=1$, which leads to values of $n_{k}$ between $10$ and $203$, with a mean of $84.2$, and the results are qualitatively similar to those in Table 1. The one notable difference to Table 1 is the better performance of the 2-way model, which is now comparable to the 3-way model in terms of the bias of the imputation estimator and outperforms both the traditional gravity model and the 1-way model in this respect. This confirms our earlier supposition that the endogeneity of the treatment explains the poor performance of the 2-way model in Table 1.
@@ 6 Conclusion
We find that the 3-way gravity model is difficult to beat when the purpose is to obtain consistent estimates of the effects of policy interventions. However, our results suggest that there may be moderate gains, especially in precision, from allowing pair characteristics with time variation to enter the model flexibly. The only area where machine learning methods can outperform the 3-way model is the prediction of individual trade flows. Given that the evaluation of policy interventions is arguably the most important reason for the use of the gravity equation in applied international trade research, we see this result as reassuring, in the sense that it further cements the role of the 3-way gravity model as the state-of-the-art specification for gravity estimation. Future research will hopefully adopt our methodological approach and investigate whether our results hold in different datasets and using different machine learning estimators.